OpenStudy (anonymous):

how can i take the derivative of this function ? f(x) =1/2x ln(x4), NOTE: this is all one thing together the two are being multiplied which is the confusing part. i can easily take the derivative of 1/2x by itself but the combination is difficult. do i do u substitutioon?

OpenStudy (anonymous):

oh and its ln(x^4)

OpenStudy (anonymous):

is it like this? \[\frac{1}{2}x\ln(x^4)\]

OpenStudy (anonymous):

its a little hard to tell whats in the denominator and what isnt.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

because i am supposed to take the derivative and then use that to find the equation of the tangent line.

OpenStudy (anonymous):

so you are going to use a combination of the product rule and chain rule. the 1/2 doesnt really matter, it will go to the front anyways. So let: \[f(x) = x, g(x) = \ln(x^4)\] then using the product rule we get: \[f'(x) = 1, g'(x) = \frac{1}{x^4}*4x^3 = \frac{4}{x}\] \[(f(x)g(x))' = f'(x)g(x)+f(x)g'(x)\]\[(x\ln(x^4))'= \ln(x^4)+x*\frac{4}{x} = \ln(x^4) + 4\]

OpenStudy (anonymous):

So putting that 1/2 back in front we get: \[\frac{1}{2}(\ln(x^4)+4)\]

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

n then what